OpenStudy (anonymous):

Find y"in terms of x and y by implicit differentiation. (x^3)+(y^3)=4^3. This problem is really stumping me. So, now i have it down to 3x+3y(-x^2/y^2)

OpenStudy (anonymous):

Where is this coming from? "(-x^2/y^2)"

OpenStudy (anonymous):

When you solve for dy/dx

OpenStudy (anonymous):

Let's take this term-by-term. What is d/dx of x^3?

OpenStudy (anonymous):

3x^2

OpenStudy (anonymous):

What is d/dx of y^3?

OpenStudy (anonymous):

3y^2(dy/dx)

OpenStudy (anonymous):

Ok, great, and 4^3 is a constant, so the derivative of that is 0. So you should have so far. \[\large 3x^2+3y^2y'=0\]

OpenStudy (anonymous):

Solve that for y'.

OpenStudy (anonymous):

y'=-x^2/y^2

OpenStudy (anonymous):

Now you want second derivative, right? So take d/dx implicitly of that.

OpenStudy (anonymous):

You can use quotient rule, or express it as \[\large -x^2y^{-2}\] and use product rule

OpenStudy (anonymous):

So, then it would be -x^2(-2y^-3)+y^-2(-2x) correct?

OpenStudy (anonymous):

Once you have that where would you go next?

OpenStudy (anonymous):

That looks good, except you're missing the y' in ((-2y'y^-3). Simplify and substitute in the y' that you already found.

OpenStudy (anonymous):

okay so then -x^2(-2y^-3(-x/y))+(y^-2)(2x) right?

OpenStudy (anonymous):

y'=-x^2/y^2.

OpenStudy (anonymous):

Okay so then -x^2(-2y^-3(-x^2/y^2))+(y^-2)(2x)? Do you go further than this?

OpenStudy (anonymous):

You can simplify it down, but that is essentially it.

OpenStudy (anonymous):

Okay. Thank you for your help.